OpenStudy (anonymous):

how to find x in 4^x+4^(1/x)=8

OpenStudy (asnaseer):

is the 2nd term definitely \(4^{\frac{1}{x}}\) and not \(4^{-x}\)?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

have you been asked to use any specific methods? e.g. solve algebraically or numerically?

OpenStudy (anonymous):

you can solve algebraically

OpenStudy (asnaseer):

I can see one answer just by observation. See if you can also spot it if you rearrange your equation as:\[4^x+4^{\frac{1}{x}}=4+4\]

OpenStudy (anonymous):

then what will be the next step?

OpenStudy (asnaseer):

observe the equation carefully - the answer should just "jump out" :)

OpenStudy (anonymous):

on which basis rearrangement is done? can it be done to solve the equation?

OpenStudy (asnaseer):

try and equate the \(4^x\) with the first 4 on the right-hand-side. Similarly try and equate the \(4^{1/x}\) with the second 4 on the right hand side. What value of x satisfies both?

OpenStudy (anonymous):

actually is this legal process? is any other method is there?

OpenStudy (asnaseer):

:) I'm not sure if it is a "legal" mathematical proof or not, but that is the method I used. Maybe others can come up with alternatives?

OpenStudy (anonymous):

can you ask this to any of your teacher (maths)?

OpenStudy (asnaseer):

I am not in school - I do maths as a hobby so I have no teacher as such :)

OpenStudy (asnaseer):

I left school many many years ago

OpenStudy (anonymous):

on which level are you?

OpenStudy (asnaseer):

I studied Aeronautical Engineering when I was at University. I now work as a software engineer

OpenStudy (anonymous):

i want to ask next question how to prove null set is subset of every set ?

OpenStudy (asnaseer):

<--- you should ask each question separately in the list to the left. you can do this by first closing this question.

OpenStudy (anonymous):

done now can you help me?